a particle is projected with velocity whose horizontal and vertical component are U and V respectively and the particle passes through a point of projection are H and K respectively

prove 2U^2K+gH=2UVH


1
Expert's answer
2021-02-19T10:29:16-0500

Given Parameters-

Let the velocity of projection be "v" , and

angle of projection be "\theta"


"H=\text{horizontal component of velocity}=vcos\theta"

"V=\text{Vertical component of velocity}=vsin\theta"


"tan\theta=\dfrac{vsin\theta}{vcos\theta}=\dfrac{V}{U}~~~~~-(1)"


The trajectory Equation for projectile motion is given by-

"y=\text{x tan}\theta-\dfrac{gx^2}{2v^2cos^2\theta}"


Since The particle Passes through "(H,K)"


"\Rightarrow K=Htan\theta-\dfrac{gH^2}{2(vcos\theta)^2}"


Substitute the above Parameter in above equation:-

"\Rightarrow K=\dfrac{HV}{U}-\dfrac{gH^2}{2U^2}"


"\Rightarrow K=\dfrac{2UHV-gH^2}{2U^2}"


"\Rightarrow 2U^2K=2UHV-gH^2"


"\Rightarrow 2U^2K+gH^2=2UHV"


Hence proved.